OpenStudy (anonymous):

The equation of a line is 3y = ax + 1 where a ≠ 0 is a constant. Given that the line has a gradient of 7/5, what is the value of a?

OpenStudy (anonymous):

a/3=7/5 a=21/5

OpenStudy (anonymous):

Thankyou VERY much

OpenStudy (anonymous):

\[3y=\frac{7}{5}+1\]\[y=\frac{21}{5}x+\frac{1}{3}\] so a =21/5